OpenStudy (anonymous):

Let f(x) be a function such that f '(x)=x(ln|x|)^2. Find the open interval(s) in which is f(x) is concave down in interval notation.

ganeshie8 (ganeshie8):

what do you notice about the tangents to the curve in blue region ?

OpenStudy (anonymous):

The tangent lines first went up and then went down

ganeshie8 (ganeshie8):

Right, more importantly notice that the "slope" of the tangent is decreasing as x is increased

ganeshie8 (ganeshie8):

The tangent keeps falling to the right in the blue region

ganeshie8 (ganeshie8):

Since f'(x) gives the slope of the tangent, can we say that f'(x) must be decreasing in the blue region ?

ganeshie8 (ganeshie8):

f(x) is concave down implies f'(x) is decreasing

ganeshie8 (ganeshie8):

f'(x) is decreasing implies f''(x) is less than 0

ganeshie8 (ganeshie8):

you're given f'(x) take the derivative and see in what intervals it is less than 0

ganeshie8 (ganeshie8):

f '(x)=x(ln|x|)^2 f''(x) = ?

OpenStudy (anonymous):

f''(x)= ln|x|(ln|x|+2) ?

ganeshie8 (ganeshie8):

Yes, solve ln|x|(ln|x|+2) < 0

OpenStudy (anonymous):

I'm not sure how to solve that, but here's my answer: ln|x|<-2

ganeshie8 (ganeshie8):

http://www.wolframalpha.com/input/?i=solve++ln |x|%28ln|x|%2B2%29+%3C+0+over+reals

ganeshie8 (ganeshie8):

AB < 0 can be split into two cases : 1) A < 0 and B > 0 2) A > 0 and B < 0

OpenStudy (anonymous):

How not sure what you mean?

OpenStudy (anonymous):

@ganeshie8 how do you find the open intervals with ln|x|<-2??

myininaya (myininaya):

do you know how to solve? \[\ln|x| \cdot (\ln|x|+2)=0 \]

OpenStudy (anonymous):

Not really..

myininaya (myininaya):

If A*B=0 then A=0 or B=0 or both=0

myininaya (myininaya):

do you know how to solve ln|x|=0?

myininaya (myininaya):

and solve also ln|x|+2=0?

myininaya (myininaya):

\[\log_a(x)=y \implies a^y=x \\ \text{ assuming } a \in (0,1 ) \cup (1,\infty),x >0\]

myininaya (myininaya):

\[\ln(x)=\log_e(x)\]

myininaya (myininaya):

have you solved the above equations I asked about?

OpenStudy (anonymous):

I don't quite get it.... e^y=x Answer: 1, -1??

myininaya (myininaya):

\[\ln|x|=0 \implies x=1 \text{ or } x=-1 \]

myininaya (myininaya):

\[\ln|x|=-2 \implies x=? \text{ or } x=?\]

myininaya (myininaya):

if the absolute value is throwing you write it like this assume x>0 then ln|x|=ln(x) solve ln(x)=0 and also solve ln(x)=-2 then assume x<0 then ln|x|=ln(-x) solve ln(-x)=0 and solve ln(-x)=-2

myininaya (myininaya):

\[\ln(x)=0 \implies x=e^{0} \implies x=1 \\ \ln(-x)=0 \implies -x=e^{0} \implies -x=1 \implies x=-1\]

myininaya (myininaya):

you can solve the other equation in a very similar way

myininaya (myininaya):

\[\ln|x|=-2 \text{ gives us two equations } \ln(x)=-2 \text{ or } \ln(-x)=-2 \] can you solve these two equations?

OpenStudy (anonymous):

I think it got it: Answer: 1/e^2 and -1/e^2 ?

myininaya (myininaya):

looks great this should be the correct order on the number line of when our f'' is 0